the length of a hall is 18 meter and its width is 13.5 meter, find the least no. of square tiles,each of side 25 cm , required to cover the floor of the hall leaving a margin of width 1.5 meter all around.

Answer by Theo(13342)

i'll convert everything to centimeters so we can be dealing with the same measures all around.

1 meter is equal to 100 centimeters, so:

length = 18 meters * 100 = 1800 centimeters.
width = 13.5 meters * 100 = 1350 centimeters.

you want to make a border around this that is 1.5 meters wide.
1.5 meters * 100 is equal to 150 centimeters.

yout total length needs to be reduced by 2 * 150 = 300 centimeters.
your total width needs to be reduced by 2 * 150 = 300 centimeters.

1800 - 300 = 1500 centimeters.
1350 - 300 = 1050 centimeters.

the total area that you need to cover is 1500 * 1050 = 1575000 square centimeters.

your tiles each have a side of 25 centimeters.

you will need 1500/25 tiles to cover the length.
you will need 1050/25 tiles to cover the width.
you will then need to multiply the number of tiles required for the length times the number of tiles required for the width to get the total number of tiles required.

1500/25 = 60
1050/25 = 42

the number of tiles required is 60 * 42 = 2520.

since the area of each tile is 25 * 25 = 625 square centimeters, then 2520 * 625 = 1575000.

this is exactly the same as the whole tiled area so your calculations are good.